My age is half my mum's age, my dad is 5 years older than my mum, the total of our ages is 100. Calculate our ages
Answer by mananth(16946)

mum's age =x
my age = (1/2)x
Dad's age = (x+5)
Add All
x+(x/2) +(x+5)=100
2x+(x/2)+5=100
multiply by 2
4x+x+10=200
5x=190
x=38.............. mum's age
my age 19 years
dad's age = 43 years